Eclipse Engineering provides services of structural engineering. They just opened a new branch in Portland, OR. One of their products is called Structural Insulated Panels, or SIPs. It is a type of foam insulation that replaces typical wall and roof framing for residential or commercial buildings. Clients are billed a fixed fee for each project based on square footage of the building, square footage of the panel area, and number of panels. Depending on how long a project takes to complete, there usually ends up being additional cost not billed to the client (fee burned). Eclipse engineering is using simple linear regression to model their costs. We will investigate the data to find the best model that can be used to predict what the fee burned will be for a particular project! Click Assignment:Submit one excel file per group and clearly label the number with each answer: ata File. 1. Calculate the correlation between each explanatory variable and the response variable (fee burned). 2 a. Perform a test for significant correlation at 5% between Panel Area and Fee burned, what is your right-tailed p- value and conclusion? b. Perform a test for significant correlation at 5% between # Panels and Fee burned, what is your right-tailed p- value and conclusion? c. Without performing the test, what do you think the results will be between Fee Burned and Floor Area? 3. Create a scatter plot for each explanatory variable and show the linear equation on each of the 3 scatter plots. (Hint: you will have to rearrange your data to create the scatter plot to list x first, then y for each.) 4. Remove the one obvious outlier you can see in the Panel Area. (Your scatter plots should update automatically.) 5. Pull 3 simple linear regression models (with outlier removed) and determine which is the best model. 6. Pull 2 more models. One with all three explanatory variables and one with some combination of 2 of the variables. ldentify your best model. 7. Write out the sample multiple regression equation and use it to predict the fee burned for taking on a project with panel area-5,000, floor area 2500, and panels -11
File Home Insert Draw Page Layout Formulas t view View Help Tell me what you want to do 四share -Com AutoSum A 多, Wrap Text General Copy Paste Conditional Format as Cell Insert Delete Format Sort &Find & .Format Painter Formatting Table StylesClear Filter Select Clipboard Font Alignment Styles Cells Editing 1The test for significant correlation: Determines whether the relationship implied by the correlation coefficient is real or due to chance Do not reject, means the correlation is zero (no correlation) Reject, means the correlation is significantly different from zero (there is significant correlation) 5 Data: Number of pairs, n 7 Alpha 8 Correlation Coefficient, r -CORREL(array1,array2) The correlation coefficient indicates direction and strength. It ranges between 1 and -1. The covariance only reveals the direction of the linear relationship between two variables. 10 Test Statistic: 11 Degrees of Freedom, df 12 Test Statistic, tf 13 Two-tailed critical value- 14 Two Tailed P-value 15 Left-tailed critical value 16 Left-tailed P-Value 17 Right-tailed critical value 18 Right-tailed P-value 19 20 21 2 n-2 , ANUM! NUM! #NUM #NUMl #NUM! 赮UM! T.INV.2t(B7, B11) T.DIST.2T(B12,B11) -T.INVfB7B11) -T. DIST(B12,B11,TRUE) -T.INVf1-B7,B11) -T.DIST.RT(B12B11) ' | 23 Fee Burned DataCorrelation test Ready
File Home Inser Draw Page Layout Formulas Data Review View Help Tell me what you want to do Share P Commer Σ AutoSum . A Copy Format Painter Paste conditional Fo matas tyles. Insert Delete Format B 1 u . . 오. .- --E t Merge & Center . s-% , +58 Fort& Selec& Clear. Formatting Table" ▼ Filter- Select" Font Alignment Number Editing 0 1 Fee Burne Panel Are Floor Area # Panels 2 1759.92 22225 67401204 12 1. Calculate the correlation between each explanatory variable and the response variable (fee burned 10 2 a. Perform a test for significant correlation at 5% between Panel Area and Fee burned, what is your right-tailed p-value and conclusion? 23 b, perform a test for significant correlation at 5% between # Panels and Fee burned. What is your right-tailed p-value and conclusion? 20 C. Without performing the test, what do you think the results will be between Fee Burned and Floor Area? 5670 5 1732.5 4478 6 1662.5 4303 2493 1012 90 1804 1296 6 3 Create a scatter plot for each explanatory variable and show the linear equation on each of the 3 scatter plots 10 (Hint: you will have to rearrange your data to create the scatter plot to list x first, then y for each.) 9 4 Remove the one obvious outlier you can see in the Panel Area. (Your scatter plots should update automatically 18 5 Pull 3 simple linear regression models (with outlier removed) and determine which is the best model. 10 Pull 2 more models. One with all three explanatory variables and one with some combination of 2 of the variables. Identify your best model. 9 7 Write othe sample multiple regression equation and use it to predict the fee burned for taking on a project with panel area 5,000, floor area 2500, and panels 11. 113814459 10 2350 51108 2491209 4168 12 130 987 2270 1160 14 15 16 17 18 19 20 857.5 928 4481 3350 1610 24 Fee Burned Data Correlation test |
Answer #1

1.between free burned(y) and panel area (X1)



X Values
∑ = 69066
Mean = 4062.706
∑(X - Mx)2 = SSx = 147711085.529

Y Values
∑ = 18234.42
Mean = 1072.613
∑(Y - My)2 = SSy = 9061658.29

X and Y Combined
N = 17
∑(X - Mx)(Y - My) = 14711481.545

R Calculation
r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy))

r = 14711481.545 / √((147711085.529)(9061658.29)) = 0.4021

Meta Numerics (cross-check)
r = 0.4021

Key

X: X Values
Y: Y Values
Mx: Mean of X Values
My: Mean of Y Values
X - Mx & Y - My: Deviation scores
(X - Mx)2 & (Y - My)2: Deviation Squared
(X - Mx)(Y - My): Product of Deviation Scores

The value of R is 0.4021.

Although technically a positive correlation, the relationship between your variables is weak (nb. the nearer the value is to zero, the weaker the relationship).

between free burned (y) and floor area(X2)


X Values
∑ = 25491
Mean = 1499.471
∑(X - Mx)2 = SSx = 7017316.235

Y Values
∑ = 18234.42
Mean = 1072.613
∑(Y - My)2 = SSy = 9061658.29

X and Y Combined
N = 17
∑(X - Mx)(Y - My) = 6201663.956

R Calculation
r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy))

r = 6201663.956 / √((7017316.235)(9061658.29)) = 0.7777

Meta Numerics (cross-check)
r = 0.7777

Key

X: X Values
Y: Y Values
Mx: Mean of X Values
My: Mean of Y Values
X - Mx & Y - My: Deviation scores
(X - Mx)2 & (Y - My)2: Deviation Squared
(X - Mx)(Y - My): Product of Deviation Scores

The value of R is 0.7777.

This is a strong positive correlation, which means that high X variable scores go with high Y variable scores (and vice versa).

between free burned (y) and # panel(X3)

X Values
∑ = 199
Mean = 11.706
∑(X - Mx)2 = SSx = 341.529

Y Values
∑ = 18234.42
Mean = 1072.613
∑(Y - My)2 = SSy = 9061658.29

X and Y Combined
N = 17
∑(X - Mx)(Y - My) = 33631.065

R Calculation
r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy))

r = 33631.065 / √((341.529)(9061658.29)) = 0.6045

Meta Numerics (cross-check)
r = 0.6045

Key

X: X Values
Y: Y Values
Mx: Mean of X Values
My: Mean of Y Values
X - Mx & Y - My: Deviation scores
(X - Mx)2 & (Y - My)2: Deviation Squared
(X - Mx)(Y - My): Product of Deviation Scores

The value of R is 0.6045.

This is a moderate positive correlation, which means there is a tendency for high X variable scores go with high Y variable scores (and vice versa).

2

a) significnce test between panel area(X1) and free burned (y)

r = 0.4021

n = 17

t test = r*square root of [(n-2)/( 1- r^2)]

t = 1.70 so, right tail P-Value is .054883.
The result is not significant at p < .05. so there is no strong correlation between panel area and free burned

b)significnce test between #panel(X3) and free burned (y)

r = 0.6045

n= 17

t =2.93 right tail P-Value is .005173.

The result is significant at p < .05 means there is a strong correlation between # panel and free burned

c) r = 0.7777 we can say that there is a strong correlation between floor area and free burned at 5 % significance because of same sample size and high r value.

3) between panel area and free burned- linear equation

y 0.0996x + 667.98 scatter plot 01617 3500 2500 1000 500 0 2000 4000 6000 8000 10000 12000 14000 16000 panel area

linear equation : y= 0.0996*x +667.98

between floor area and free burned- linear equation

scatter plot y 0.8838x-252.57 R2 0.6048 3500 2500 2000 u 1500 1000 500 500 1000 1500 2000 2500 floor area

linear equation : y = 0.8838*x - 252.57

between# panel and free burned- linear equation

scatter plot y98.472x-80.088 R2 0.3655 3500 2500 2000 u 1500 1000 500 10 15 20 25 # panel

linear equation : y = 98472*x -80.088

4) after removal the outlier

14459-panel area 1138- free burned

y 0.4256x 383.86 scatter plot P250676586 3500 2500 E 2000 1500 1000 500 0 1000 2000 3000 4000 5000 6000 7000 8000 panel area

y = 0.4256*x -383.86

5) between panel area and free burned after removal outlier

Y = -383.8617 + 0.4256 X1

Source

DF

Sum of Squares Mean Square F Statistic P-value
Regression
(between ŷi andyi bar)

1

5953159.9054

5953159.9054

26.8510

0.0001391

Residual
(between yi and ŷi)

14

3103955.7000

221711.1214

Total(between yi andyi bar)

15

9057115.6054

603807.7070

Coeff

SE t-stat lower t0.025(14) upper t0.975(14)

Stand Coeff

p-value

VIF

intercept -383.8617 304.0027 -1.2627 -1035.8825 268.1592 0.000 0.2273
slope 0.4256 0.08212 5.1818 0.2494 0.6017 0.8107 0.0001391 1.0000

Y and X relationship
R square (R2) equals 0.6573. It means that the predictors (Xi) explain 65.7% of the variance of Y.
Adjusted R square equals 0.6328.
The coefficient of correlation (R) equals 0.8107. It means that there is a very strong direct relationship between the predicted data (ŷ) and the observed data (y).

Goodness of fit
Overall regression: right-tailed, F(1,14) = 26.8510, p-value = 0.0001391. Since p-value < α (0.05), we reject the H0.
The linear regression model, Y = b0+ b1X1provides a better fit

All the independent variables (Xi) are significant.

The Y-intercept : two-tailed, T = -1.2627, p-value = 0.2273. Hence intercept is not significantly different from zero. It is still most likely recommended not to force b to be zero.

between floor area and free burned - there is no outlier

Y = -252.5679 + 0.8838 X1

Source

DF

Sum of Squares Mean Square F Statistic P-value
Regression
(between ŷi andyi bar)

1

5480818.3840

5480818.3840

22.9589

0.0002378

Residual
(between yi and ŷi)

15

3580839.9055

238722.6604

Total(between yi andyibar)

16

9061658.2896

566353.6431

Coeff

SE t-stat lower t0.025(15) upper t0.975(15)

Stand Coeff

p-value

VIF

intercept -252.5679 300.8844 -0.8394 -893.8878 388.7521 0.000 0.4144
slope 0.8838 0.1844 4.7915 0.4906 1.2769 0.7777 0.0002378 1.0000

Y and X relationship
R square (R2) equals 0.6048. It means that the predictors (Xi) explain 60.5% of the variance of Y.
Adjusted R square equals 0.5785.
The coefficient of correlation (R) equals 0.7777. It means that there is a strong direct relationship between the predicted data (ŷ) and the observed data (y).

Goodness of fit
Overall regression: right-tailed, F(1,15) = 22.9589, p-value = 0.0002378. Since p-value < α (0.05), we reject the H0.
The linear regression model, Y = b0+ b1X1 provides a better fit t

All the independent variables (Xi) are significant.

The Y-intercept : two-tailed, T = -0.8394, p-value = 0.4144. Hence intercept is not significantly different from zero. It is still most likely recommended not to force b to be zero.

between # panel and free burned - there is no outlier

y = -80.0880 + 98.4719 X1

Source

DF

Sum of Squares Mean Square F Statistic P-value
Regression
(between ŷi andyibar)

1

3311716.2806

3311716.2806

8.6393

0.01015

Residual
(between yi and ŷi)

15

5749942.0089

383329.4673

Total(between yi andyibar)

16

9061658.2896

566353.6431

Coeff

SE t-stat lower t0.025(15) upper t0.975(15)

Stand Coeff

p-value

VIF

intercept -80.0880 419.9374 -0.1907 -975.1634 814.9873 0.000 0.8513
slope 98.4719 33.5021 2.9393 27.0639 169.8800 0.6045 0.01015 1.0000

Y and X relationship
R square (R2) equals 0.3655. It means that the predictors (Xi) explain 36.5% of the variance of Y.
Adjusted R square equals 0.3232.
The coefficient of correlation (R) equals 0.6045. It means that there is a strong direct relationship between the predicted data (ŷ) and the observed data (y).

Goodness of fit
Overall regression: right-tailed, F(1,15) = 8.6393, p-value = 0.01015. Since p-value < α (0.05), we reject the H0.
The linear regression model, Y = b0+ b1X1, provides a better fit.

All the independent variables (Xi) are significant.

The Y-intercept : two-tailed, T = -0.1907, p-value = 0.8513. Hence intercept is not significantly different from zero. It is still most likely recommended not to force b to be zero.

so, the best model is panel area nad free burned relationship.

6)

with three explanatory variable

y = -740.43 + 0.0454749*X1 + 0.659653*X2 + 54.6017X3

Variable Parameter S.D. T-STAT
H0: parameter = 0
2-tail p-value 1-tail p-value
(Intercept) -740.4 336.6 -2.2000e+00 0.04651 0.02326
X1 +0.04548 0.03758 +1.2100e+00 0.2478 0.1239
X2 +0.6596 0.1887 +3.4960e+00 0.003946 0.001973
X3 +54.6 26.01 +2.1000e+00 0.05586 0.02793


Multiple Linear Regression - Regression Statistics
Multiple R 0.8532
R-squared 0.7279
Adjusted R-squared 0.6651
F-TEST (value) 11.59
F-TEST (DF numerator) 3
F-TEST (DF denominator) 13
p-value 0.0005619
Multiple Linear Regression - Residual Statistics
Residual Standard Deviation 435.5
Sum Squared Residuals 2.465e+06

WITH two explanatory variable - floor area and # panel

y = -650.516 + 0.722435X2 + 54.6614V3X3

Variable Parameter S.D. T-STAT
H0: parameter = 0
2-tail p-value 1-tail p-value
(Intercept) -650.5 333.7 -1.9500e+00 0.07155 0.03578
X2 +0.7224 0.1844 +3.9170e+00 0.001548 0.0007738
X3 +54.66 26.43 +2.0680e+00 0.05766 0.02883


Multiple Linear Regression - Regression Statistics
Multiple R 0.835
R-squared 0.6973
Adjusted R-squared 0.654
F-TEST (value) 16.12
F-TEST (DF numerator) 2
F-TEST (DF denominator) 14
p-value 0.0002329
Multiple Linear Regression - Residual Statistics
Residual Standard Deviation 442.6
Sum Squared Residuals 2.743e+06

Based on F test and slope test of two models we can say that the second model is best ( two variable) because 69 % explained by this two variable if I include third variable help to explained extra 3 %.

7) y = -740.43 + 0.0454749*X1 + 0.659653*X2 + 54.6017X3

x1 =5000

x2= 2500

x3 =11

y = -740.43 + 0.0454749*5000 + 0.659653*2500 + 54.6017*11

y = 1736.6957